### v2

Abstract
We provide Dutch-book arguments against two forms of misspecified Bayesian learning. An agent progressively learns about a state and is offered a bet after every new discovery. We say the agent is Dutch-booked when they are willing to accept all bets, but their payoff is negative under each state either ex-post, or in expectation given the objective conditional probabilities of the discoveries (i.e., the correct data-generating process, DGP). Respectively, an agent cannot be Dutch-booked if and only if they update their beliefs with Bayes rule either from the previous belief, even using misspecified likelihood functions, or from one lexicographic prior, using the correct data-generating process. Under a large population interpretation of the DGP, this means that a population can suffer aggregate losses under all states when different individuals update their beliefs from different (lexicographic) priors, or using misspecified likelihoods. Thus, the Dutch-book argument offers a general characterization of the perils of misspecification.
Keywords: Dutch book, Bayesian updating, lexicographic beliefs, model misspecification
1 Introduction
The classical Dutch-book argument (Ramsey 1931, De Finetti 1937) is a cornerstone of probability theory. It says that an agent (he) may accept a set of bets that return an aggregate loss under every state (i.e., a Dutch book) if his preferences cannot be rationalized by one probabilistic belief. Later (Lehman 1955, Kemeny 1955), the converse was also proven: an agent cannot be Dutch-booked if he evaluates all bets with the same probabilistic belief about the state.
In this Dutch-book argument there is no time dimension. But the agent may progressively receive information about a fundamental state, and the bookmaker (she) could offer one bet after every piece of news. In this paper, we consider the following learning environment. There are a set of decision-relevant fundamental states and a set of contingencies, partially ordered in time, in which the agent makes choices. The fundamental states represent different (probabilistic) DGPs for contingencies. Thus, the contingencies have objective probabilities conditional on the fundamental states. Then, a question arises: Should the beliefs at counterfactual contingencies be somehow coherent to guarantee sound decision-making? We answer this question with a two-fold Dutch-book argument. [^1] Epistemologists have studied this problem in a framework where the agent learns finer and finer partitions of the state space (see Rescorla, 2022 for a recent survey). Under the assumption that every state is deemed possible at the outset, Teller (1973) and Lewis (1999) have shown that an agent cannot be Dutch-booked if and only if he updates beliefs by conditioning. Rescorla (2022) extends the argument to the presence of subjectively impossible states. While in the partitional framework each state can only give rise to one learning path, and thus counterfactual contingencies can only be consistent with different states, in the framework of this paper a state can result into different paths, and thus be consistent with different counterfactual contingencies.
We call deterministic Dutch book a system of bets on the fundamental state, one for each contingency, so that the cumulative payoff of the agent is non-positive under all fundamental states and paths, and negative under some fundamental state and path. More generally, we call Dutch book a system of bets that yields a non-positive expected payoff under every fundamental state, and a negative expected payoff under some fundamental state, where the expectations are computed with the probabilities of the contingencies as determined by the fundamental state under consideration. An agent is (deterministically) Dutch-booked when she is willing to accept all the bets in a (resp. deterministic) Dutch book. [^2] We use the following tie-breaking rule: the agent does not accept bets with zero subjective expected payoff.
In other words, a deterministic Dutch book guarantees that under every possible DGP the agents surely receives a negative payoff, and possibly gets a strictly negative payoff. Instead, a Dutch book requires that the agent’s expected payoff given the probabilities over contingencies induced by every DGP is nonpositive, and strictly negative for some DGP.
The notion of deterministic Dutch book extends to this paper’s framework the notion of “diachronic Dutch book” employed by the arguments for conditioning. With this, we show that an agent cannot be deterministically Dutch-booked if and only if his beliefs are forward-consistent, i.e., the agent updates beliefs with Bayes rule whenever possible. Forward-consistency does not require to use the correct conditional probabilities of the contingencies, so the agent cannot be deterministically Dutch-booked even if they are misspecified about the relative probabilities in the likelihood function, i.e., they may update on the basis of a wrong likelihood function that links the fundamental states and the contingencies. It only rules out the more radical form of misspecification where a state is subjectively believed possible given some contingency that is objectively impossible under that state. Moreover, forward-consistency is equivalent to Bayes-plausibility (Kamenica and Gentzkow, 2011), i.e., the requirement that each current belief is a convex combination of the beliefs at the next contingencies. Thus, Bayes-plausibility is the coherence rule among beliefs at counterfactual contingencies that characterizes an agent who cannot be deterministically Dutch-booked.
The necessity of forward-consistency to avoid Dutch-booking does not follow from the existing results that establish the necessity of conditioning in the partitional framework (see Rescorla, 2022 for the most general case), because of the multiplicity of learning paths that are compatible with the same state. Thus, given a violation of forward-consistency between a contingency and its successors, the bookmaker needs to find an acceptable bet for the first contingency such that the opposite bet will be accepted in each of the next contingencies (despite the potentially different posteriors), and not just in one as in the partitional framework. [^3] Another way of seeing this difficulty is that, in this paper, the bookmaker is constrained to offer bets on one dimension of the “grand state space” that also includes the learning paths. The solution to this problem will be an application of Farkas lemma.
Leveraging the separation between subjective (over fundamental states) and objective (given the fundamental state) uncertainty, the general notion of Dutch book aggregates payoffs across counterfactual contingencies according to the objective probability over them prescribed by the fundamental state. Such weaker form of Dutch book is important for two orders of reason.
First, a population of agents making choices in different contingencies will realize the aggregate payoff under the true state. So, if the population can be Dutch-booked in objective expected terms, their combined choices will select sure aggregate losses and forfeit sure aggregate wins, suggesting an evolutionary disadvantage. Alternatively, a single agent may perish in the long run after being exposed to i.i.d. occurrences of the same dynamic decision problem.
Second, the currently accepted interpretation of Dutch-book arguments is not the literal one, where a bookmaker can truly design a bet for each contingency, but rather that the agent suffers from some form of irrationality, which the Dutch-book argument brings to light.
We illustrate how general Dutch books work with an example.
Example 1
Larry wants to go out and watch the New Year’s fireworks, but the municipality does not disclose their location anymore to avoid crowding. There are three possible locations: main square, marina, and central park. There are three bus lines on his street: the red line goes to the square, the blue line to the marina, and the green line to the park. The buses always arrive in this order, at constant time intervals. Larry is late, so his only chance to see the fireworks is taking the first bus that arrives, regardless of its color, and hoping it goes to the right location. Suppose the blue line comes. When Larry arrives at the marina, he discovers whether the fireworks are there or not. If they are not, Larry regrets not moving to the bus stop a little earlier, which meant catching the red line: “I bet the fireworks are at the square!” Overhearing this, a bookmaker approaches Larry and proposes a bet on the fireworks’ location. Larry accepts a bet that yields $9$ if the fireworks are at the square and $-10$ if the fireworks are at the park. By the same token, at the park, he would accept a bet that yields $9$ for marina and $-10$ for square, and at the square, he would accept a bet that yields $9$ for park and $-10$ for marina. Thus, no matter the fireworks’ location, Larry ends up with equal probability in each of the other two locations, and hence he suffers an expected loss from this betting behavior.
In the example, the bus schedule determines the objective conditional probabilities of the three contingencies of interest. Instead, there are no objective probabilities of the fireworks’ location, and Larry does not have a subjective prior either; after all, he cannot decide which bus to take, so he does not need one. However, to deal with the bookmaker, he needs to form a belief after discovering that the fireworks are not in the location he arrived to. What’s wrong with the beliefs he forms? The location he deems more likely is different in each of the three contingencies. This form of incoherence among beliefs can arise precisely from the lack of a prior. Thus, the Dutch-book argument provides a motivation for having a prior, even when there is no decision to make before the arrival of information. More generally, the argument provides a justification for having coherent beliefs that can be rationalized with a prior, although the agent may not truly have one, but just reason in a coherent way after different discoveries. In the population interpretation, having a prior translates into the agents having the same prior. Thus, the Dutch book arguments suggests a disadvantage for “divided” populations, in which different agents have different (prior) beliefs, in that the agents with different priors can be exploited in different situation. Such divided population naturally emerges in a repeated setting if the agents do not share their experiences, thus suggesting an important role for the aggregation of feedback.
Another possible reason for Larry’s incoherent beliefs is that he (has a prior, but) perceives an inexistent correlation between bus arrivals and location of the fireworks. This is a classic form of model misspecification about the likelihood function connecting fundamental states and contingencies. In fact, we show that any form of model misspecification about the likelihood function allows Dutch-booking. This is perhaps surprising: since there is no “bad prior” over fundamental states (because there is no objective one), it is hard to conceive in what sense updating the prior with wrong likelihoods could generate a “bad posterior” over fundamental states. The Dutch-book argument offers a precise sense in which model misspecification is bad for decision-making.
In this simple example, having one prior and updating it by conditioning suffices to avoid Dutch-booking. But this is not true in general. We determine the coherence rule among beliefs that is necessary and sufficient to avoid Dutch-booking in objective expected terms. The coherence rule is easy to interpret; it is based on the idea that differences in the subjective odds ratios of states across contingencies shall only reflect differences in the objective odds ratios of the contingencies given the states. In the example, if Larry deems square twice more likely than park when he is at the marina and marina twice more likely than square when he is at the park, to satisfy the coherence rule he should deem marina four times more likely than park when he is at the square.
In turn, we show that this coherence rule characterizes a Completely Consistent Belief System: an array of beliefs (one for each contingency) that can be derived from a Lexicographic Conditional Probability System (Blume et al., 1991), a list of theories in a priority order. Specifically, each belief is derived with Bayes rule from the first theory that is consistent with the contingency. So, the LCPS plays the role of lexicographic prior, while the CCBS specifies the posterior beliefs in different contingencies. The fact that a completely consistent belief rather than simply conditioning from a unique prior is needed will be illustrated in Section 6 with an example of an agent who is subject to correlation neglect.
Related literature
In the context of sequential games, Siniscalchi (2022) introduces the notion of structural rationality, which is based on the use of CCBSs over the opponents’ strategies. [^4] Interestingly, structural rationality is motivated by the inability of sequential rationality to capture coherent conditional preferences that can be derived by updating a prior preference relation. CCBSs guarantee that all conditional preferences can be derived from a prior lexicographic preference relation. In other words, they guarantee a stronger form of Bayesianism which might be called “Bayesian updating of preferences” compared to Bayesian updating of beliefs. In Siniscalchi (2022), as well as Battigalli et al. (2023), a CCBS is defined as an array of beliefs, one for each information set, that is consistent with a Complete Conditional Probability System: an array of probability measures, one for each subset of opponents’ strategies, disciplined by the chain rule of probability. [^5] Conditional probability systems were first introduced by Renyi (1955). CCPSs were introduced in the analysis of games by Myerson (1986). In turn, a CCPS can be derived from a LCPS by conditioning. In a sequential game, a player progressively rules out opponents’ strategies by observing their moves, that is, learning occurs in the partitional fashion. So, Bayes rule boils down to conditioning, and hence the definition of CCBS in Siniscalchi (2022) and Battigalli et al. (2023) is a special case of the definition of this paper. Thus, the Dutch-book argument for complete consistency implies that the standard notion of conditional probability system defined on the collection of observable events falls short of the level of coherence that shelters from Dutch-booking in objective expected terms. The working paper of Siniscalchi (2022) also characterizes CCBSs with a generalization of the chain rule that applies between non-nested conditional events. In the Appendix, we provide a direct proof of the equivalence between the generalized chain rule and our coherence rule in sequential games.
Pettigrew (2023) considers an agent with a credence at time $t$ and a vector of possible credences at $t^{\prime}>t$, and shows that the agent can be Dutch-booked if and only if his credence at $t$ is not in the convex hull of the credences at $t^{\prime}$ (which he calls “General Reflection Principle”). This result yields our equivalence between Bayes-plausibility and the impossibility of deterministic Dutch-booking for the special case of a tree with depth $1$ and discoveries that do not rule out any state. [^6] When some discoveries objectively rule out some states, statement II.a of Theorem 1 in Pettigrew (2023) fails. From his theorem, Pettigrew also derives an argument for conditioning in the partitional framework. In our richer framework, we can derive our more general argument for forward consistency, uncovering the limited possibility of achieving a deterministic Dutch-booking against likelihood misspecification.
Chen et al. (2015) show that the posteriors of different agents, who update a common prior based on different discoveries, are common knowledge if and only if they cannot be Dutch-booked in aggregate. [^7] See also Lehrer and Samet (2011) for an alternate characterization of “agreeing to agree” between only two agents. Their notion of Dutch book only takes the realized discoveries and sums up the payouts under each state. This amounts to assuming that the bookmaker can choose the bets based on the entire set of realized discoveries. So, when the agents have different (thus not commonly known) posteriors, the bookmaker can exploit an informational advantage over each single agent, who is unaware of the other agents’ discoveries. In other words, the bookmaker can make profit from arbitrage. This is ruled out by our notion of Dutch book, in which each possible discovery is associated with a bet, independently of what other discoveries (using the population interpretation) may have been made by other agents. This reconciles the possibility of Dutch-booking Bayesian agents in Chen et al. (2015) with the impossibility results of this paper.
There are several related contribution in the literature on misspecified Bayesian learning, who is excellently surveyed in Bohren and Hauser (2024a). Lanzani (2025) characterizes a behavior rule that prevents misspecification to damage the agent beyond his maxmin payoff preserving consistency when correctly specified. Bohren and Hauser (2024b) characterizes when a departure from Bayesian updating (e.g., underinference from signals) can be rationalized as the consequence of Bayesian updating within a misspecified model. Molavi (2024) shows that beliefs are consistent with Bayesian updating (from a possibly misspecified model) if and only if the mean posterior is absolutely continuous with respect to the prior. This finding generalizes the earlier Shmaya and Yariv (2016) to the case in which the subjective beliefs of the decision maker may not have the same support of the true data-generating process. Cerreia-Vioglio et al. (2025) show that in coordination games misspecification can be extremely damaging even for misspecification concerned agents. Compared to this literature, we provide a novel reason for why some misspecification should or should not be expected to persist based on their susceptibility to Dutch books.
Finally, our results are similar in spirit to those in the no-trade literature, the closest being in Morris (1994). Morris considers agents who trade claims over commodities that are contingent on a payoff-relevant state and on the realization of a signal. In the case of two traders, a single money commodity and risk neutrality, Morris’ model coincides with ours given a tree of depth $1$. Assuming that the traders have no incentive to exchange claims that are contingent on the state only (which in our setting boils down to having the same prior), Theorem 2.1 in Morris (1994) shows that there exists trade that gives a positive expected payoff to each trader if and only if they disagree on the likelihood function mapping states into probabilities over signals. Our main theorem further shows (in a more general learning framework) that such trade can take a very specific form, whereby one trader has a positive expected payoff conditional on each state, and the other has a positive expected payoff conditional on each signal. If the first trader (the bookmaker) is using the correct likelihood function, then the second trader (our agent) gets an objectively negative expected payoff conditional on each state, and thus is Dutch-booked. Conversely, we show that the existence of such trade guarantees that the disagreement between the two traders must be about the likelihood function, not (just) the prior.
The paper is organized as follows. Section 1 introduces the framework. Section 2 illustrates belief systems. Section 3 introduces the novel notions of Dutch book. Section 4 states and proves the Dutch-book theorem for forward-consistency. Section 5 is devoted to the Dutch-book theorem for complete consistency. In the Appendix, we formalize the special case of sequential games and the connection with Siniscalchi (2022).
2 Framework
There is a finite set $\mathcal{S}$ of possible (fundamental) states. We describe the possible learning processes through an arborescence (i.e., a set of trees) of contingencies $\mathcal{H}$. With this, we are not assuming the existence of just one initial contingency in which the agent forms beliefs before learning anything about the state. A learning path is a path of contingencies from an “initial” contingency that is not preceded by any other contingency (i.e., the root of a tree), to a “terminal” contingency that is not followed by any other contingency (i.e., a leaf of that tree). Let $\mathcal{L}$ denote the set of possible learning paths and $L(h)$ the set of paths that go through a contingency $h$, and $H(l)$ the set of contingencies through which learning path $l$ goes. Each state $s\in\mathcal{S}$ induces an objective distribution $\eta^{s}\in\Delta(\mathcal{L})$ over learning paths, and thus states can be interpreted as different possible data-generating processes. Let $L(s)=\mathrm{supp}\eta^{s}$. For each $h\in\mathcal{H}$, let $p(h|s)=\sum_{l\in L(h)}\eta^{s}(l)$ denote the probability of reaching contingency $h$ under state $s$. Let $S(h)$ denote the set of states $s$ that are consistent with contingency $h$, that is, $p(h|s)>0$. Without loss of generality, we assume that every contingency is consistent with some state. An important special case is the one in which the probability of a contingency $h$ is the same across all the consistent states $s$, that is, $p(h|s)=p(h|s^{\prime})=:p^{h}$ for all $h\in\mathcal{H}$ and $s,s^{\prime}\in S(h)$. We will refer to this as the “independence case”.
To illustrate the framework, we formalize the example of the Introduction.
Example 2
There are three possible states, corresponding to the location of the fireworks: $\mathcal{S=}\left\{sq,ma,pa\right\}$. There are six possible contingencies, each with a different set of consistent states, so we label each contingency $h$ with the only state or with the initials of the two states in $S(h)$:
=$\mathcal{H=}\left\{sq,ma,pa,sm,mp,ps\right\}.$
For instance, if Larry takes the bus that goes to the square, he ends up in contingency $sq$ if the fireworks are there, and in contingency $mp$ otherwise. Note that all the contingencies are unordered, therefore contingencies and learning paths coincide. The arrival process of the buses is independent of the state. Thus, we are in the independence case: for each $h\in\left\{sm,mp,ps\right\}$ and $s,s^{\prime}\in S(h)$, $p(h|s)$ and $p(h|s^{\prime})$ are equal, because they coincide with the probability of the bus that leads to contingency $h$. In particular, since the three lines arrive with equal frequencies, each state $s\in\mathcal{S}$ induces a uniform distribution $\eta^{s}$ over the three learning paths/contingencies $h$ such that $s\in S(h)$. So, $p(h|s)=p(h|s^{\prime})=1/3$. $\triangle$
3 Belief systems
We assume that, in each contingency $h$, the agent has a probabilistic belief $\mu(\cdot|h)\in\Delta(\mathcal{S})$ such that $\mu(S(h)|h)=1$ — we call $(\mu(\cdot|h))_{h\in\mathcal{H}}$ belief system. This maintained assumption, besides requiring probabilistic sophistication at any point in time, rules out the most severe forms of model misspecification in which an agent assigns positive probability to states that have become objectively impossible.
We say that a belief system is forward-consistent if the agent’s beliefs are consistent with Bayesian updating from his previous belief using some (conditional) likelihood function, whenever possible. Observe that such conditional likelihood function is not required to be the objectively correct one.
Definition 1
A belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is forward-consistent if, for every non-terminal $h\in\mathcal{H}$, calling $h^{1},...,h^{n}$ its immediate successors, there exist conditional probability distributions
$\left(\widetilde{p}(\cdot|h,s)\right)_{s\in S(h)}\in\left(\Delta(\left\{h^{1},...,h^{n}\right\})\right)^{\left|S(h)\right|}$
such that, for each $m\in\left\{1,...,n\right\}$, if $\widetilde{p}(h^{m}|h,s)>0$ for some $s\in\mathrm{supp}\mu(\cdot|h)\subseteq S(h)$, then $\mu(\cdot|h^{m})$ can be derived with Bayes rule from $\mu(\cdot|h)$ using $(\widetilde{p}(h^{m}|h,s))_{s\in S(h)}$.
It is well-known that beliefs are consistent with this form of updating if and only if the “prior belief” (i.e., the belief at $h$) is a convex combination of the “posterior beliefs” (i.e., the beliefs at $h^{1},...,h^{n}$). (The Appendix contains a formal proof of this fact.)
Definition 2
A belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is Bayes-plausible if, for every non-terminal $h\in\mathcal{H}$, calling $h^{1},...,h^{n}$ its immediate successors, there exist $\alpha^{1},...,\alpha^{n}\in\left[0,1\right]$ such that $\alpha^{1}+...+\alpha^{n}=1$ and
$\sum_{m=1}^{n}\alpha^{m}\mu(\cdot|h^{m})=\mu(\cdot|h).$
As we will show, Bayes-plausibility is the minimal requirement of coherence among beliefs that shelters from deterministic Dutch-booking.
We will define a Completely Consistent Belief System as a belief system that can be derived from a Lexicographic Conditional Probability System.
Definition 3 (Blume et al., 1991)
A list of probability measures $\bar{\mu}=(\mu^{1},...,\mu^{n})$ over $\mathcal{S}$ is a Lexicographic Conditional Probability System if for every $s\in\mathcal{S}$, there exists exactly one $m\in\left\{1,...,n\right\}$ such that $\mu^{m}(s)>0$.
Definition 3 embodies a “full-support” requirement: every state is given positive probability at some level of the LCPS. In this way, from every LCPS, one can always derive a belief system as follows.
Definition 4
A belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is completely consistent if there exists a LCPS $\bar{\mu}=(\mu^{1},...,\mu^{n})$ such that, for each $h\in\mathcal{H}$, $\mu(\cdot|h)$ can be derived with Bayes rule using $\left(\eta^{s}\right)_{s\in\mathcal{S}}$ from the earliest $\mu^{m}$ such that $\mu^{m}(S(h))>0$, i.e., $m=\min\left\{k:\mu^{k}(S(h))>0\right\}$ and
$\mu(s|h)=\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}\in S}p(h|s^{\prime})\mu^{m}(s^{\prime})}\qquad\forall s\in S(h)\text{.}$
[…]
We will characterize CCBSs in terms of coherence among the subjective relative probabilities of states in different contingencies. However, given that the objective probabilities of contingencies may differ across states, these comparisons are meaningful only after discounting the probability attributed to a state in a contingency by the probability that the state induces the contingency.
Definition 5
Given a belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$, a contingency $h$, and a pair of states $s,s^{\prime}\in S(h)$ such that $\mu(s|h)+\mu(s^{\prime}|h)>0$, the (possibly infinite) discounted odds ratio between $s$ and $s^{\prime}$ at $h$ is
$o(s,s^{\prime}|h)=\frac{\mu(s|h)}{p(h|s)}\frac{p(h|s^{\prime})}{\mu(s^{\prime}|h)}.$
A generalized odds ratio between $s$ and $s^{\prime}$ is the product of a finite concatenation of discounted odds ratios
$o(s,s^{1}|h^{1})\cdot o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-1},s^{\prime}|h^{n})$
such that, if one discounted odds ratio is 00, no discounted odds ratio is infinite.
A discounted odds ratio compares the probabilities attributed to two states in the same contingency. A generalized odds ratio compares the probabilities of two states, $s$ and $s^{\prime}$, in two different contingencies, $h^{1}$ and $h^{n}$, via a chain of comparisons with other states ($s$ with $s^{1}$, $s^{1}$ with $s^{2}$, …, until $s^{n-1}$ with $s^{\prime}$) in different contingencies ($h^{1}$ for $s,s^{1}$, $h^{2}$ for $s^{1},s^{2}$, and so on). [^8] The formal definition allows contingencies and states to repeat, because this will be convenient for the proofs. In the independence case, discounted odds ratios boil down to (standard) odds ratios.
The strong rule of coherence that shelters from Dutch-booking in objective expected terms is the equivalence of all generalized odds ratios between any two states.
Definition 6
We say that a belief system is completely coherent when, for every $s,s^{\prime}\in\mathcal{S}$, the generalized odds ratios between $s$ and $s^{\prime}$ are all identical.
We illustrate the notions of belief system and generalized odds ratio through the example from the Introduction.
Example 3
For each state $s$, a belief system $\mu=\left(\mu(\cdot|h)\right)_{h\in\mathcal{H}}$ must assign probability $1$ to $s$ in the contingency labelled with $s$. Also, in this example, every belief system is trivially forward-consistent as every path consists of just one (terminal) contingency.
To obtain a CCBS, let us fix a LCPS $\bar{\mu}=(\mu^{1},...,\mu^{n})$. Suppose that $\mu^{1}$ assigns positive probability to at least $2$ states. Then, $\mu^{1}(S(h))>0$ for each $h\in\left\{sm,mp,ps\right\}$. Because each contingency has the same probability under the two consistent states, we are in the independent case: Bayes rule boils down to the chain rule, and discounted odds ratios boil down to odds ratios. So, for each $h\in\left\{sm,mp,ps\right\}$, writing $S(h)=\left\{s,s^{\prime}\right\}$, we have
$\mu(s|h)=\frac{\mu^{1}(s)}{\mu^{1}(s)+\mu^{1}(s^{\prime})}.$
If instead $\mu^{1}$ assigns probability $1$ to a state $\bar{s}$, we obtain $\mu(\bar{s}|h)=1$ whenever $\bar{s}\in S(h)$, and the belief in the contingency $h$ such that $S(h)=\mathcal{S}\backslash\left\{\bar{s}\right\}=\left\{s,s^{\prime}\right\}$ is given by
$\mu(s|h)=\frac{\mu^{2}(s)}{\mu^{2}(s)+\mu^{2}(s^{\prime})}.$
Given any two states $s,s^{\prime}$, one can compute an odds ratio between $s$ and $s^{\prime}$ in the only contingency $h$ such that $S(h)=\left\{s,s^{\prime}\right\}$; call it $o(s,s^{\prime})$. Then, one can also compute a generalized odds ratio between $s$ and $s^{\prime}$ as $o(s,s^{\prime\prime})\cdot o(s^{\prime\prime},s^{\prime})$ (s′′{$s^{\prime\prime}\notin\left\{s,s^{\prime}\right\}$), unless one is zero and the other infinite. For instance, when $\mu^{1}(ma)\in(0,1)$, we can compute
$o(sq,ma)\cdot o(ma,pa)$ $=$ $\frac{\mu(sq|sm)}{\mu(ma|sm)}\cdot\frac{\mu(ma|mp)}{\mu(pa|mp)}=\frac{\mu^{1}(sq)}{\mu^{1}(ma)}\cdot\frac{\mu^{1}(ma)}{\mu^{1}(pa)}$
$=$ $\frac{\mu^{1}(sq)}{\mu^{1}(pa)}=\frac{\mu(sq|ps)}{\mu(pa|ps)}=o(sq,pa),$
which can be infinite if $\mu^{1}(pa)=0$ and thus $\mu^{1}(sq)>0$, or zero, vice versa. Note the equality between the odds ratio and the generalized odds ratio.
An alternative belief system $\bar{\mu}$, which rationalizes Larry’s betting behavior, is the following:
$\bar{\mu}(sq|ps)=\bar{\mu}(ma|sm)=\bar{\mu}(pa|mp)=3/4\text{.}$
In this case, the odds ratio and the generalized odds ratio between any two states differ. For instance, in contingency $ps$ we have $o(sq,pa)=3$, but in contingency $sm$ we have $o(sq,ma)=1/3$ and in contingency $mp$ we have $o(ma,pa)=1/3$, therefore we can compute a generalized odds ratio between $sq$ and $pa$ of $1/9$ instead of $3$triangle
4 Dutch books and Dutch-booking
Traditionally, a Dutch book is a set of bets such that, no matter the realization of the state, the cumulative payoff from the bets is never positive and sometimes negative for the gambler (thus never negative and sometimes positive for the bookmaker). An agent is Dutch-booked if he accepts all the bets in a Dutch book. Our agent, however, would never accept a classical Dutch book if all the bets are proposed at the same time, as he has a well-defined probabilistic belief at any point in time. We now investigate whether, instead, a system of gambles proposed in different contingencies can induce the same kind of state-wise losses. [^9] Our agent risk neutrality and probabilistic sophistication at every contingency imply that, when he is willing to accept a bet, he is also willing to accept the sum of the bets, and thus we can assume without loss of generality that our agent is proposed just one bet in each contingency.
Definition 7
A gamble is a map $\gamma:\mathcal{S}\rightarrow\mathbb{R}$ that specifies, for each $s\in\mathcal{S}$, the gain or loss $\gamma(s)$ for the agent in case of $s$.
We say that $\gamma$ is a gamble on $\widetilde{S}\subseteq\mathcal{S}$ if $\gamma(s)=0$ for every $s\in\mathcal{S}\backslash\widetilde{S}$.
A system of gambles is an array $g=(g(\cdot|h))_{h\in\mathcal{H}}$ where, for each $h\in\mathcal{H}$, $g(\cdot|h)$ is a gamble on $S(h)$. [^10] The results would not change if the bets in a Dutch book were allowed to yield non-zero payoffs also under objectively impossible states.
What does it mean, for a system of gambles, to be a “Dutch book”? An intuitive notion is that the bookmaker obtains a never negative and sometimes positive profit under every state and path. We call this notion “deterministic Dutch book”. But the existence of objective conditional probabilities of paths allows to introduce also a more permissive notion of Dutch book, whereby the bookmaker obtains a non-negative/positive profit conditional on each funamental state only in objective expected terms.
Definition 8
A system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$ is a Dutch book if for every $s\in\mathcal{S}$,
$\sum_{h\in\mathcal{H}}p(h|s)g(s|h)\leq 0,$ (1)
and for some $s\in\mathcal{S}$,
$\sum_{h\in\mathcal{H}}p(h|s)g(s|h)<0.$ (2)
A system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$ is a deterministic Dutch book if for every $s\in\mathcal{S}$ and $l\in L\left(s\right)$
$\sum_{h\in H(l)}g(s|h)\leq 0,$ (3)
and for some $s\in\mathcal{S}$ and $l\in L\left(s\right)$,
$\sum_{h\in H(l)}g(s|h)<0.$ (4)
Note that a deterministic Dutch book does not depend on the objective probabilities of paths, hence it is a Dutch book no matter such probabilities. Moreover, Dutch books, deterministic or not, do not depend on the objective probabilities of the states.
Since each gamble is proposed to the agent in a specific contingency, the acceptance of the gamble will depend on the contingent belief.
Definition 9
Fix a probability measure $\nu\in\Delta(\mathcal{S})$ and a gamble $\gamma$. We say that the agent is willing to accept $\gamma$ given $\nu$ if
$\sum_{s\in\mathcal{S}}\nu(s)\gamma(s)>0\text{,}$
or $\gamma(s)=0$ for every $s\in\mathcal{S}$.
We say that an agent with belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is willing to accept a system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$ if, for every $h\in\mathcal{H}$, he is willing to accept $g(\cdot|h)$ given $\mu(\cdot|h)$.
We say that the agent is (deterministically) Dutch-booked if he is willing to accept a (deterministic) Dutch book.
Definition 9 captures the following tie-breaking rule: our agent declines gambles that yield zero subjective expected payoff (except for the trivial “null” gamble, as the bookmaker is free to offer no gamble). This rules out in a simple way the uninteresting case in which the agent can be Dutch-booked by just charging negative payoffs for the states he subjectively deems impossible.
To conclude this section, we illustrate Dutch-booking in the running example.
Example 4
Suppose that Larry has the belief system $\bar{\mu}$ constructed before. The bookmaker constructs the following system of bets:
$g(sq|ps)$ $=$ $g(ma|sm)=g(pa|mp)=9;$
$g(pa|ps)$ $=$ $g(sq|sm)=g(ma|mp)=-10;$
$g(s|h)$ $=$ 0 $0\text{ \ \ \qquad\qquad\qquad\qquad\qquad\qquad\qquad}\forall h\in\left\{sq,ma,pa\right\}.$
Under every state $s$, we have
$\sum_{h\in\mathcal{H}}p(h|s)g(s|h)=\frac{1}{3}(9-10+0)=-\frac{1}{3}<0,$
i.e., Larry has an objective expected payoff of $-1/3$. Therefore, the system is a Dutch book. There remains to show that Larry is willing to accept it. We have
$\sum_{s\in\mathcal{S}}\bar{\mu}(s|h)g(s|h)=\frac{3}{4}\cdot 9+\frac{1}{4}(-10)+0\cdot 0>0\quad\forall h\in\left\{sm,mp,ps\right\}.$
Therefore, Larry is Dutch-booked. Note that deterministic Dutch-booking is impossible in this example, as every path consists of just one contingency, in which the agent has a well-defined probabilistic belief. $\triangle$
5 The Dutch-book theorem for forward-consistency
[…]
Theorem 1
Consider an agent with belief system $\mu=(\mu(\cdot|h))_{h\in\mathcal{H}}$. The following are equivalent:
  1. 1.
    $\mu$ is Bayes-plausible;
  2. 2.
    $\mu$ is forward-consistent;
  3. 3.
    the agent cannot be deterministically Dutch-booked.
Note that forward consistency does not depend on the models the agent uses for updating. In this sense, misspecification about the relative probabilities of the likelihood function is immune to deterministic Dutch-booking. However, that is not the case for the stronger form of misspecification where at some contingency $h$ some objectively impossible state $s\notin S\left(h\right)$ is deemed possible by the decision maker. This is because, by assigning an arbitrarily high positive payout $g(s|h)>>0$ to the agent in $s$ given $h$, the bookmaker can obtain a profit under some state $s^{\prime}\in S\left(h\right)$ that is objectively possible.
Proof of Theorem 1.
$\mathbf{1\Rightarrow 2)}$
This is a known fact, however we provide formal proof in the Appendix.
$\mathbf{2\Rightarrow 3)}$
[…]
$\mathbf{3\Rightarrow 1)}$
We prove the counterpositive. Suppose that $\mu$ is not Bayes-plausible. Thus, there exists a non-terminal contingency $h$ such that $\mu(\cdot|h)$ is not a convex combination of the beliefs at the next contingencies $h^{1},...,h^{n}$. Let $A\in\mathbb{R}^{\left|\mathcal{S}\right|\times n}$ be the matrix of subjective probabilities $\mu(s|h^{j})$, and let $b\in\mathbb{R}^{\left|\mathcal{S}\right|}$ be the vector of subjective probabilities $\mu(s|h)$. So, for every transposed vector $x=(x^{1},...,x^{n})\in\left[0,1\right]^{n}$ such that $\sum_{m=1}^{n}x^{m}=1$, $Ax\not=b$. Moreover, for every $x\in\left[0,\infty\right)^{n}$ such that $\sum_{m=1}^{n}x^{m}\not=1$, $Ax\not=b$, because $Ax=b$ implies
$x^{1}+...+x^{n}$ $=$
$\sum_{j=1}^{n}\left(\sum_{s}\mu(s|h^{j})\right)x^{j}$ $=$
$\sum_{s}\sum_{j=1}^{n}x^{j}\mu(s|h^{j})$ $=$
$\sum_{s}\mu(s|h)$ $=$ $1$
Thus, there is no $x\geq 0$ such that $Ax=b$. But then, by Farkas lemma, there exists $y\in\mathbb{R}^{\left|\mathcal{S}\right|}$ such that
$A^{T}y\geq 0\text{ and }b^{T}y<0.$
For some $\varepsilon>0$, by $-b^{T}y>0$ the gamble with vector of prizes $-y-2\overrightarrow{\varepsilon}$ is acceptable at $h$, and by $A^{T}y\geq 0$ the gamble with vector of prizes $y+\overrightarrow{\varepsilon}$ is acceptable at each $h^{j}$. Complete the system with gambles with zero prizes for all states at every other contingency. We obtain a deterministic Dutch book: if one of the paths that go through $h$ realizes, and those includes at least one $l\in L(s)$ for some $s\in S(h)$, the profit of the bookmaker is $\varepsilon>0$ (satisfying 4), otherwise it is 00 (satisfying 3). $\blacksquare$
6 The Dutch-book theorem for complete consistency
The Dutch-book theorem for complete consistency generalizes the insights from the example: incongruities among generalized odds ratios can be exploited for Dutch-booking and can only be displayed by a belief system that is not completely consistent. The theorem also shows the opposite implications: when there is no incongruity between generalized odds ratios, the agent’s belief system is completely consistent and he cannot be Dutch-booked. Thus, coherence among generalized odds ratios fully characterizes complete consistency and impossibility of Dutch-booking.
Theorem 2
Fix a belief system $\mu$. The following are equivalent.
  1. 1.
    $\mu$ is completely coherent;
  2. 2.
    $\mu$ is completely consistent;
  3. 3.
    the agent cannot be Dutch-booked.
Proof. The acronyms DOR and GOR will stand for discounted/generalized odds ratio.
$\mathbf{1\Rightarrow 2)}$
Note preliminarly that, by Condition 1, if a state $s$ has a DOR with another state $s^{\prime}$, all the DORs between $s$ and $s^{\prime}$ take the same value; [^12] Recall that an “indeterminate DOR” betwen $s$ and $s^{\prime}$ is just not considered a DOR, and likewise for GOR. We say that $s$ has a DOR with $s^{\prime}$ when there exists $h\in\mathcal{H}$ such that $s,s^{\prime}\in S(h)$ and $\mu(s|h)+\mu(s^{\prime}|h)>0$, the DOR between $s$ and $s^{\prime}$ being $\mu(s|h)/\mu(s^{\prime}|h)$. Similarly, we say that $s$ has a GOR with $s^{\prime}$ when it is possible to compute at least one GOR between $s$ and $s^{\prime}$. call it $o(s,s^{\prime})$.
Let $P^{1}$ be the set of all $s\in\mathcal{S}$ that have no zero GOR with any $s^{\prime}$.
Let $P^{2}$ be the set of all $s\in\mathcal{S}\backslash P^{1}$ that have no zero GOR with any $s^{\prime}\in\mathcal{S}\backslash P^{1}$.
Let $P^{3}$ be the set of all $s\in\mathcal{S}\backslash\left(P^{1}\cup P^{2}\right)$ that have no zero GOR with any $s^{\prime}\in\mathcal{S}\backslash\left(P^{1}\cup P^{2}\right)$.
And so on.
Claim 1: there exists $n\geq 1$ such that $(P^{1},...,P^{n})$ is a partition of $\mathcal{S}$, with non-empty elements.
Proof: Let $n$ be the smallest non-negative integer such that $P^{n+1}=\emptyset$; it exists because $\mathcal{S}$ is finite and the sets are disjoint by construction. There remains to show that $\cap_{m=1}^{n}P^{m}=\mathcal{S}$ (which also implies that $n\geq 1$). Let $\widetilde{S}=\mathcal{S}\backslash\left(\cap_{m=1}^{n}P^{m}\right)$ and suppose by contraposition that $\widetilde{S}\not=\emptyset$. The fact that $P^{n+1}=\emptyset$ implies that every $s\in\widetilde{S}$ has a zero GOR with some $s^{\prime}\in\widetilde{S}$. Consider a directed graph with nodes $\widetilde{S}$ and arrows that represent a zero GOR between the start and the end node. Since every node is the start of an arrow, the graph has a cycle. So, the product of all the (zero) GORs along the cycle is a zero GOR between $s^{1}$ and itself. Inverting all the DORs in this GOR, we obtain an infinite GOR between $s^{1}$ and itself. This violates the assumption that all the GORs between $s^{1}$ and itself are identical by Condition 1. $\square$
Claim 2: there exists a LCPS $\bar{\mu}=(\nu^{1},...,\nu^{n})$ such that, for each $m=1,...,n$, (a) $\mathrm{supp}\nu^{m}=P^{m}$ and (b), for every pair $\left(s,s^{\prime}\right)\in P^{m}\times P^{m}$ which has a DOR $o(s^{\prime},s)$, $\nu^{m}(s^{\prime})/\nu^{m}(s)=o(s^{\prime},s)$.
Proof: For each $m=1,...,n$, we will construct a probability measure $\nu^{m}$ that satisfies (a) and (b); then, $(\nu^{1},...,\nu^{n})$ is an LCPS because, by Claim 1, $(P^{1},...,P^{n})$ partitions $\mathcal{S}$.
Fix $m=1,...,n$ and construct a graph with nodes $P^{m}$ and connections between nodes that have a DOR. For every connected component $\widetilde{P}\subseteq P^{m}$, pick a state $s$ arbitrarily, let $\widetilde{\nu}(s)=1$, and progressively let $\widetilde{\nu}(s^{\prime\prime})=\widetilde{\nu}(s^{\prime})\cdot o(s^{\prime\prime},s^{\prime})$ for every $s^{\prime\prime}\in\widetilde{P}$, where $o(s^{\prime\prime},s^{\prime})$ is a DOR between $s^{\prime\prime}$ and some $s^{\prime}\in\widetilde{P}$ for which $\widetilde{\nu}(s^{\prime})$ was already defined; by definition of $P^{m}$, $o(s^{\prime\prime},s^{\prime})$ is not zero and not infinite (otherwise $o(s^{\prime},s^{\prime\prime})$ would be zero). So, we can obtain a probability measure $\nu^{m}$ with support $P^{m}$ by rescaling $\widetilde{\nu}$.
To show that $\nu^{m}$ satisfies the desiderata, fix $s,s^{\prime}\in P^{m}$ that have a DOR $o(s^{\prime},s)$, thus $s$ and $s^{\prime}$ are in the same connected component. Then, the graph has a path $(\bar{s}^{1},...,\bar{s}^{T})$ with $\bar{s}^{1}=s$ and $\bar{s}^{T}=s^{\prime\prime}$ such that, for each $t=2,...,T$, $\widetilde{\nu}(\bar{s}^{t})=\widetilde{\nu}(\bar{s}^{t-1})\cdot o(\bar{s}^{t},\bar{s}^{t-1})$, where $o(\bar{s}^{t},\bar{s}^{t-1})\in\mathbb{R}^{++}$ is a DOR between $\bar{s}^{t}$ and $\bar{s}^{t-1}$. (Note that, if $\widetilde{\nu}(\bar{s}^{t-1})$ was defined as $\widetilde{\nu}(\bar{s}^{t})\cdot o(\bar{s}^{t-1},\bar{s}^{t})$ and not vice versa, there is also a DOR $o(\bar{s}^{t},\bar{s}^{t-1})=1/o(\bar{s}^{t-1},\bar{s}^{t})$ between $\bar{s}^{t}$ and $\bar{s}^{t-1}$.) Therefore
$\frac{\nu^{m}(s^{\prime})}{\nu^{m}(s)}=\frac{\widetilde{\nu}(s^{\prime})}{\widetilde{\nu}(s)}\equiv\frac{\widetilde{\nu}(\bar{s}^{T})}{\widetilde{\nu}(\bar{s}^{1})}=o(\bar{s}^{T},\bar{s}^{T-1})\cdot\frac{\widetilde{\nu}(\bar{s}^{T-1})}{\widetilde{\nu}(\bar{s}^{1})}=...=\prod\limits_{t=2}^{T}o(\bar{s}^{t},\bar{s}^{t-1}).$
This is a GOR between $s^{\prime}$ and $s$, and it is equal to $o(s^{\prime},s)$ by Condition 1. $\square$
Now fix $h\in\mathcal{H}$. Let $m$ be the smallest $k$ such that $\nu^{k}(S(h))>0$. We show that $\mu(\cdot|h)$ can be derived from $\nu^{m}$ with Bayes rule. Let $\hat{\nu}^{m}$ be the probability measure derived from $\nu^{m}$ with Bayes rule given $h$ (and $\left(\eta^{s}\right)_{s\in\mathcal{S}}$). By property (a) of $\bar{\mu}$, $\mathrm{supp}\nu^{m}=P^{m}$, thus $\mathrm{supp}\hat{\nu}^{m}=P^{m}\cap S(h)=:\hat{S}$. Moreover, for all $s,s^{\prime}\in\hat{S}$,
$\frac{\hat{\nu}^{m}(s)}{\hat{\nu}^{m}(s^{\prime})}=\frac{p(h|s)\nu^{m}(s)}{p(h|s^{\prime})\nu^{m}(s^{\prime})}.$
To show that $\mu(\cdot|h)$ and $\hat{\nu}^{m}$ coincide, it is enough to show that $\mathrm{supp}\mu(\cdot|h)=\hat{S}$ and $\mu(s|h)/\mu(s^{\prime}|h)=\hat{\nu}^{m}(s)/\hat{\nu}^{m}(s^{\prime})$ for all $s,s^{\prime}\in\hat{S}$.
First we show that $\mathrm{supp}\mu(\cdot|h)=P^{m}\cap S(h)$.
Fix $s\in\mathrm{supp}\mu(\cdot|h)$. Thus, $s\in S(h)$. Hence, for every $k<m$, $s\not\in P^{k}$, as $\mathrm{supp}\nu^{k}=P^{k}$ but $\nu^{k}(S(h))=0$. Now we show that, for every $k>m$, $s\not\in P^{k}$ either, so that $s\in P^{m}$. Suppose by contradiction that $s\in P^{k}$. As $\mathrm{supp}\nu^{m}=P^{m}$ and $\nu^{m}(S(h))>0$, there exists $\bar{s}\in S(h)\cap P^{m}$ such that $\nu^{m}(\bar{s})>0$. Since $s\in\mathrm{supp}\mu(\cdot|h)$, $s$ has a non-zero DOR with every $s^{\prime}\in S(h)$, so in particular with $\bar{s}$. Since $s\not\in P^{1}\cup...\cup P^{m}$, $s$ has a zero GOR with some $s^{\prime}\in P^{m}\cup...\cup P^{n}$. The two things combined impy that $\bar{s}$ has a zero GOR with $s^{\prime}$ too. But this contradicts $\bar{s}\in P^{m}$.
Now fix $s\not\in\mathrm{supp}\mu(\cdot|h)$. If $s\not\in S(h)$, we are done. If $s\in S(h)$, $s\not\in P^{m}$, because $s$ has a zero DOR with every $s^{\prime}\in\mathrm{supp}\mu(\cdot|h)$ and we have just shown that $\mathrm{supp}\mu(\cdot|h)\subseteq P^{m}$.
To conclude the proof, fix a pair $s,s^{\prime}\in\mathrm{supp}\mu(\cdot|h)$ and write
$\frac{\mu(s|h)}{\mu(s^{\prime}|h)}=\frac{p(h|s)}{p(h|s^{\prime})}\frac{\mu(s|h)}{p(h|s)}\frac{p(h|s^{\prime})}{\mu(s^{\prime}|h)}=\frac{p(h|s)}{p(h|s^{\prime})}o(s,s^{\prime})=\frac{p(h|s)\nu^{m}(s)}{p(h|s^{\prime})\nu^{m}(s^{\prime})},$
where $o(s,s^{\prime})$ is the DOR between $s$ and $s^{\prime}$ at $h$, and the last equality follows from property (b) of $\nu^{m}$.
$\mathbf{2\Rightarrow 3)}$
Let $\bar{\mu}=(\mu^{1},...,\mu^{n})$ be a LCPS from which the belief system is derived. For each $m=1,...,n$, let $S^{m}$ denote the support of $\mu^{m}$, and let $H^{m}$denote the set of all $h\in\mathcal{H}$ such that $\mu(\cdot|h)$ is derived from $\mu^{m}$.
Fix a system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$. Suppose that $i$ is willing to accept $g$. We show that $g$ is not a Dutch book. Suppose that $g$satisfies condition (1) of a Dutch book, so that,
$\forall s\in\mathcal{S},\text{ \ \ }\sum_{h\in\mathcal{H}}p(h|s)g(s|h)\leq 0.$ (5)
We are going to show that $g(s|h)=0$ for all $s\in\mathcal{S}$ and $h\in\mathcal{H}$, so that $g$ does not satisfy condition (2) of a Dutch book.
The proof is by induction on $m$. Fix $m=1,...,n$ and suppose by way of induction that for every $k<m$, $g(s|h)=0$ for all $s\in\mathcal{S}$ and $h\in H^{k}$. (This is vacuously true for $m=1$.) Fix $s\in S^{m}$. For every $k>m$ and $h\in H^{k}$, $p(h|s)=0$, otherwise $\mu(\cdot|h)$ could be derived from $\mu^{m}$. By this and the induction hypothesis, inequality (5) yields
$\sum_{h\in H^{m}}p(h|s)g(s|h)\leq 0.$
So, considering that $\mu^{m}(s^{\prime})=0$ for every $s^{\prime}\not\in S^{m}$, we can write
$\sum_{s\in S}\mu^{m}(s)\sum_{h\in H^{m}}p(h|s)g(s|h)\leq 0\text{.}$ (6)
Since $\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})>0$ for every $h\in H^{m}$, we can rewrite (6) as
$\sum_{h\in H^{m}}\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)\leq 0.$ (7)
For each $h\in H^{m}$, since the agent is willing to accept $g(\cdot|h)$, we have
$\sum_{s\in\mathcal{S}}\mu(s|h)g(s|h)\geq 0.$ (8)
Since $\mu(\cdot|h)$ can be derived from $\mu^{m}$ with Bayes rule, we can rewrite (8) as
$\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)\geq 0.$ (9)
Equations (9) and (7) imply
$\forall h\in H^{m},\text{ \ \ }\sum_{s\in S}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)=0,$
that is,
$\sum_{s\in S}\mu(s|h)g(s|h)=0.$
This means that the agent has no strict incentive to accept $g(\cdot|h)$, therefore acceptance implies $g(s|h)=0$ for all $s\in\mathcal{S}$.
$\mathbf{3\Rightarrow 1)}$
The proof that an agent with two different GORs between a state $s$ and a state $s^{\prime}$ can be Dutch-booked resembles the classical Dutch-book arguments of De Finetti and Ramsey. Therefore, the proof is deferred to the Appendix. $\blacksquare$
The next example illustrates the theorem.
Example 5
This is a variation of the famous Monty Hall problem, in which the initially selected door is determined at random and not revealed to the decision maker. There are three doors. A car is placed behind one of the doors, and a goat is placed behind each of the other two doors. The state represents the door with the car: $\mathcal{S}=\left\{A,B,C\right\}$. Then, one door, call it $d$, is drawn from a uniform distribution (in the Monty Hall problem, this is the door assigned to the agent). Finally, another door behind which there is a goat, call it $h\not=d$, is opened; it is the first non-car door preceding $d$ in an $a$-$b$-$c$ circle. So, for instance, conditional on $s=B$, door $a$ is opened with probability $2/3$ (the probability that $d\in\left\{b,c\right\}$), and door $c$ is opened with probability $1/3$ (the probability that $d=a$). The agent only observes which door is opened, so the possible contingencies are: =$\mathcal{H}=\left\{a,b,c\right\}$. Overall, conditional on each $s\in\mathcal{S}$, the probabilities of each $h\in\mathcal{H}$ are:
$p(h|s)$ $h=a$ $h=b$ $h=c$
$s=A$ 00 $\frac{1}{3}$ $\frac{2}{3}$
$s=B$ $\frac{2}{3}$ 00 $\frac{1}{3}$
$s=C$ $\frac{1}{3}$ $\frac{2}{3}$ 00
We next consider two possible belief systems for the agent.
First, suppose that, at every $h\in\mathcal{H}$, the agent assigns equal probability to each of the two states in $S(h)$. Then, the agent can be Dutch-booked (in objective expected terms). To see this, compute the three discounted odds ratios:
$o(B,C|a)$ $=$ $\frac{\Pr(a|C)}{\Pr(a|B)}=\frac{1/3}{2/3}=\frac{1}{2};$
$o(C,A|b)$ $=$ $\frac{\Pr(b|A)}{\Pr(b|C)}=\frac{1/3}{2/3}=\frac{1}{2};$
$o(B,A|c)$ $=$ $\frac{\Pr(c|A)}{\Pr(c|B)}=\frac{2/3}{1/3}=2.$
Since
$o(B,C|a)\cdot o(C,A|b)\not=o(B,A|c),$
we have two different generalized odds ratios between $B$ and $A$, and thus by Theorem 2 the agent can be Dutch-booked. For example, suppose that, at each $h\in\mathcal{H}$, a bookmaker offers to the agent the following bets:
$g(C|a)$ $=$ 3, $3,\text{ \ }g(B|a)=-2;$
$g(A|b)$ $=$ 3, $3,\text{ \ }g(C|b)=-2;$
$g(B|c)$ $=$ 3, $2\displaystyle 3,\text{ \ }g(A|c)=-$ 2 .
Clearly, the agent accepts all bets, and the objective expected payoff of the bookmaker conditional on each $s\in\mathcal{S}$ is:
$-g(A|b)p(b|A)-g(A|c)p(c|A)$ $=$ $-g(B|c)p(c|B)-g(B|a)p(a|B)=-g(C|a)p(a|C)-g(C|b)p(b|C)$
[…]
So, the agent is harmed by the “correlation neglect” he experiences.
Second, suppose that the agent derives his beliefs as follows. He has a uniform prior, and, thinking of the correct updating in the original Monty Hall problem, believes that the car is more likely to be behind the unopened door that has not been assigned to him (i.e., not door $d$). For instance, suppose that $h=a$ and the agent correctly computes
$\Pr\left(d=b|h=a\right)$
$=$ $\frac{\Pr(h=a|d=b)\Pr(d=b)}{\Pr(h=a|d=b)\Pr(d=b)+\Pr(h=a|d=c)\Pr(d=c)}$
$=$ $\frac{\Pr(h=a|d=b,s=B)+\Pr(h=a|d=b,s=C)}{\left(\begin{array}[]{c}\Pr(h=a|d=b,s=B)+\Pr(h=a|d=b,s=C)\\ +\Pr(h=a|d=c,s=B)+\Pr(h=a|d=c,s=C)\end{array}\right)}$
$=$ $\frac{2}{2+1}=\frac{2}{3}.$
So, he has $\mu(s=B|h=a)<1/2$. However, given $h=a$, the car is more likely to be behind door $b$:
$\Pr\left(s=B|h=a\right)$
$=$ $\frac{\Pr(h=a|s=B)\Pr(s=B)}{\Pr(h=a|s=B)\Pr(s=B)+\Pr(h=a|s=C)\Pr(s=C)}$
$=$ $\frac{\Pr(s=B)}{\Pr(s=B)+\Pr(d=b)\Pr(s=C)}$
$=$ $\frac{1}{1+\frac{1}{3}}=\frac{3}{4}.$
In other words, the agent perceives the opposite correlation between the fundamental state and the observed contingency. This increases the discrepancy between generalized odds ratios, making it easier for the bookmaker to Dutch-book the agent. [^13] To be precise, without limited liability any discrepancy allows the bookmaker to extract an infinite amount of money (in expectation) from the agent. But under limited liability, the higher discrepancy allows to strictly increase the bookmaker’s expected profit.
7 Conclusion
Dutch-book arguments have already been used to motivate probabilistic sophistication and conditioning. We introduce novel Dutch-book arguments that motivate two forms of belief consistency in a learning framework a la Anscombe-Aumann, with separation between subjective uncertainty (the state) and objective uncertainty (lotteries over learning paths conditional on the state). Bayes-plausibility is necessary and sufficient to avoid being Dutch-booked in a deterministic sense, i.e., under each realized state and learning path. A stronger form of coherence, guaranteed by deriving all beliefs from a subjective lexicographic prior with the correct models, is necessary and sufficient to avoid being Dutch-booked “in objective expected terms”, i.e., under each realized state, given the corresponding lottery on the learning paths. The latter form of Dutch-booking entails sure aggregate losses across a population of agents, thus showing the perils of heterogenous priors across different individuals, and of model misspecification even in absence of an objective prior.
Appendix A Appendix
A.1 Sequential games: connection with Battigalli et al. (2023) and Siniscalchi (2020, 2022)
In the context of a sequential game with complete information and perfect recall, Battigalli et al. (2023) introduce a classification of the belief systems of a player over the opponents’ strategies. We study their classification of belief systems as a special case of the analysis of the main body of the paper. In a game, the moves of a player are independent of the moves of the co-players that are yet to be observed, and the two jointly determine the path of the game, thus the learning path. Therefore, the assumption that every contingency is reached with the same probability under every consistent state is satisfied in that context. For concreteness, we stick to the sequential games application, but the analysis extends seamlessly to situations where this assumption is satisfied, for instance in situations where the an agent does not choose what to observe but the factors that determine the observations are independent of the state itself, as in the running example of the paper.
Let $S_{-i}$ be the set of stategy profiles of the opponents of player $i$. Let $H_{i}$ be the arborescence of information sets of player $i$, endowed with the precedence relation $\prec$. Each information set $h_{i}$ corresponds to the observation of some past moves of the opponents, summarized by the event $S_{-i}(h_{i})\subseteq S_{-i}$. A belief system is an array $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ of probability measures over $S_{-i}$ such that $\mu_{i}(S_{-i}(h_{i})|h_{i})=1$ for every $h_{i}\in H_{i}$.
Taking as primitive a collection of conditioning events $\mathcal{C\subseteq}2^{S_{-i}}\backslash\left\{\emptyset\right\}$, a conditional probability system over $S_{-i}$ is an array of probability measures $\bar{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C))_{C\in\mathcal{C}}$ such that $\bar{\mu}_{i}(C|C)=1$ for every $C\in\mathcal{C}$, and for every $D\in\mathcal{C}$ with $D\subset C$, for each $E\subseteq D$,
$\bar{\mu}_{i}(E|C)=\bar{\mu}_{i}(E|D)\bar{\mu}_{i}(D|C).$
A conditional probability system is complete when it is defined with respect to the collection $\mathcal{C=}2^{S_{-i}}\backslash\left\{\emptyset\right\}$.
Definition 10 (Battigalli et al., 2023)
A belief system $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ over the opponents’ strategies is forward consistent if, for all $h_{i},h_{i}^{\prime}\in H_{i}$, if $h_{i}\prec h_{i}^{\prime}$, then
$\mu_{i}(E|h_{i})=\mu_{i}(S_{-i}(h_{i}^{\prime})|h_{i})\mu(E|h_{i}^{\prime})$
for every $E\subseteq S_{-i}(h_{i}^{\prime})$.
Analogously to the general notion of Definition 1, forward consistency corresponds to Bayesian updating of conjectures along the path of play (with the inconsequential difference that consistency is not only required between a contingency and its immediate successors). However, in this particular setting Bayesian updating boils down to the chain rule.
Definition 11
[…]
$\mu_{i}(\cdot|h_{i})=\bar{\mu}_{i}(\cdot|S_{-i}(h_{i})).$
The definition of complete consistency of Battigalli et al. (2023) uses CCPS’s to discipline beliefs at counterfactual contingencies. An information set $h$ does not carry any additional information about the strategies of the opponents besides ruling out all the strategies that do not belong to $S_{-i}(h)$; for this reason, it is appropriate to derive each belief by mere conditioning in place of Bayes rule, from the most general theory that explains the contingency.
Proposition 1
For every information set $h_{i}$, for every pair $s_{-i},s_{-i}^{\prime}\in S_{-i}(h_{i})$, and for every probability distribution over player $i$’s strategies, the probability of reaching the information set is identical under $s_{-i}$ and $s_{-i}^{\prime}$: $p(h_{i}|s_{-i})=p(h_{i}|s_{-i}^{\prime})$.
Proof. Under each $s_{-i}$, the probability of reaching $h_{i}$ is the probability of the strategies of $i$ that, with $s_{-i}$, induce a path that goes through $h_{i}$. So all we need to show is that such set of strategies of $i$ is the same under all $s_{-i}\in S_{-i}(h_{i})$. By perfect recall, all these strategies must prescribe the same moves before reaching $h_{i}$. Suppose by contradiction that there exists a strategy of $i$ that induces a path that goes through $h_{i}$ with some $s_{-i}\in S_{-i}(h_{i})$ but not with some other $s_{-i}^{\prime}\in S_{-i}(h_{i})$. But then, all other strategies of $i$ that prescribe the same moves before $h_{i}$, will have the same effect: $h_{i}$ is not reached under $s_{-i}^{\prime}$. This contradicts that $s_{-i}^{\prime}$ is consistent with $h_{i}$. $\blacksquare$
Although the use of CCPSs is appropriate in this context, one might as well model the theories from which a player derives just as a LCPS, and derive each belief by conditioning the first theory in the list that is consistent with the contingency; the two definitions of complete consistency are equivalent. [^14] The equivalence between LCPSs and CCPSs is well-known. We formalize and prove this result for the analysis to be self-contained.
Proposition 2
Fix a belief system $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$. The following are equivalent:
  1. 1.
[…]
    $\mu_{i}(\cdot|h_{i})=\bar{\mu}_{i}(\cdot|S_{-i}(h_{i}));$
  2. 2.
    there exists a LCPS $\hat{\mu}_{i}=(\mu_{i}^{1},...,\mu_{i}^{n})$ such that for each $h_{i}\in H_{i}$ and $s_{-i}\in S_{-i}(h_{i})$,
    $\mu_{i}(s_{-i}|h_{i})=\frac{\mu_{i}^{k}(s_{-i})}{\mu_{i}^{k}(S_{-i}(h_{i}))},$
    where $k$ is the smallest $m$ such that $\mu_{i}^{k}(S_{-i}(h_{i}))>0$.
Proof.
[…]
1$\Rightarrow$2) Create a decreasing chain of conditioning events $(C^{1},...,C^{n})$ as follows. Let $C^{1}=S_{-i}$. For each $k>1$, let $C^{k}$ be the union of all the $s_{-i}\in S_{-i}$ such that $\bar{\mu}_{i}(s_{-i}|C^{j})=0$ for every $j<k$. Let $n$ be the smallest $k$ such that $C^{k+1}=\emptyset$. Consider the list of measures $\hat{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C^{1}),...,\bar{\mu}_{i}(\cdot|C^{n}))$. For each $h_{i}\in H_{i}$, let $k$ be the smallest $j\leq n$ such that $\bar{\mu}_{i}(S_{-i}(h_{i})|C^{j})>0$ — it exists by $C^{n+1}=\emptyset$. By $\bar{\mu}_{i}(S_{-i}(h_{i})|C^{j})=0$ for every $j<k$, it follows that $S_{-i}(h_{i})\subseteq C^{j}$. Hence, $\mu_{i}(\cdot|h_{i})$ can be derived from $\bar{\mu}_{i}(\cdot|C^{k})$ by updating, and $\bar{\mu}_{i}(\cdot|C^{k})$ is also the first measure in $\hat{\mu}_{i}$ that gives positive probability to $S_{-i}(h_{i})$. There remains to show that $\hat{\mu}_{i}$ is a LCPS. For each $s_{-i}\in S_{-i}$, by $C^{n+1}=\emptyset$, $\bar{\mu}_{i}(s_{-i}|C^{k})>0$ for some $k\leq n$, so $\hat{\mu}_{i}$ has full joint support. Moreover, by $s_{-i}\in C^{k}$, $\bar{\mu}_{i}(s_{-i}|C^{j})=0$ for every $j<k$, and by $\bar{\mu}_{i}(s_{-i}|C^{k})>0$, $s_{-i}\not\in C^{j}$ for every $j>k$, which implies $\bar{\mu}_{i}(s_{-i}|C^{j})=0$. Hence, $\hat{\mu}_{i}$ has disjoint supports and is therefore a LCPS. $\blacksquare$
Proposition 2 implies that the main result of this paper also applies to belief consistency as it was defined for sequential games. With this, the characterization of complete consistency with the coherence rule among discounted odds ratios can be rewritten using traditional odds ratios in place of discounted odds ratios, because in this context the two coincide by Proposition 1.
Definition 12
Given a belief system $(\mu_{i}(\cdot|h))_{h\in H_{i}}$, an information set $h$, and a pair $s_{-i},s_{-i}^{\prime}\in S_{-i}(h)$, let
$o^{\ast}(s_{-i},s_{-i}^{\prime}|h)=\frac{\mu_{i}(s_{-i}|h)}{\mu_{i}(s_{-i}^{\prime}|h)}$
denote the odds ratio between $s_{-i}$ and $s_{-i}^{\prime}$ at $h$, provided that it is not indeterminate. A simple generalized odds ratio between $s_{-i}$ and $s_{-i}^{\prime}$ is the product of a finite concatenation of odds ratios
$o^{\ast}(s_{-i},s_{-i}^{1}|h^{1})\cdot o^{\ast}(s_{-i}^{1},s_{-i}^{2}|h^{2})\cdot....\cdot o^{\ast}(s_{-i}^{n-1},s_{-i}^{\prime}|h^{n})$
that is not indeterminate.
The following is a corollary of the Dutch-book theorem for complete consistency and of the last two propositions.
Corollary 1
Fix a belief system $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$. The following are equivalent:
the belief system is completely consistent;
for every $s_{-i},s_{-i}^{\prime}\in S_{-i}$, all the simple generalized odds ratios between $s_{-i}$ and $s_{-i}^{\prime}$ are identical.
Siniscalchi (2020) (the working paper of Siniscalchi 2022) comes to a similar characterization of complete consistency starting from a different idea. Consider a belief system that is isomorphic to a conditional probability system defined over the collection of observable events. Such belief system thus satisfies the chain rule between contingencies that correspond to nested observations, but the chain rule may have little bite when the observable events are not nested, and the corresponding beliefs are not disciplined by the preliminary observation of a larger event. Therefore, Siniscalchi (2020) generalizes the chain rule to apply also between overlapping, but non-nested observable events. We conclude this section by showing the convergence between Siniscalchi’s approach and mine.
Call “generalized self-odds ratio” a generalized odds ratio between $s_{-i}$ and itself.
Remark 1
All the generalized odds ratios of each pair are identical if and only if all generalized self-odds ratios are $1$.
Proof. Only if. Fix a generalized self-odds ratios. Break it into two and invert one of the two halves. We have two concatenations of odds ratios that start and end with the same element. Thus, the two corresponding generalized odds ratios are identical by assumption. Therefore, the generalized self-odds ratio is $1$.
If. Fix two concatenations of odds ratios where either no odds ratio is infinite or there is no zero. If both contain a zero, the corresponding generalized odds ratios are both zero. If they both contain an infinite, they are both infinite. If one contains an infinite and the other contains a zero, the generalized self-odds ratio obtained by inverting one of the two is zero or infinite, which violates the assumption. Otherwise, the generalized self-odds ratio (in one direction) is finite. By assumption, it is $1$. Therefore the two generalized odds ratios are identical. $\blacksquare$
Definition 13 (Siniscalchi 2020)
A belief system $(\mu^{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ is consistent if for every $(h_{i}^{1},...,h_{i}^{n})\in H_{i}$ and $E\subseteq S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{n})$,
$\mu^{i}(E|h_{i}^{1})\prod_{m=1}^{n-1}\mu^{i}(S_{-i}(h_{i}^{m})\cap S_{-i}(h_{i}^{m+1})|h_{i}^{m+1})=\mu^{i}(E|h_{i}^{n})\prod_{m=1}^{n-1}\mu^{i}(S_{-i}(h_{i}^{m})\cap S_{-i}(h_{i}^{m+1})|h_{i}^{m}).$ (11)
Theorem 3
A belief system $(\mu^{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ is consistent if and only if it is completely consistent.
Proof. We will use the characterization of complete consistency of Corollary 1.
If part. Fix $(h_{i}^{1},...,h_{i}^{n})\in H_{i}$ and $E\subseteq S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{n})$. If both sides of equation 11 are zero, consistency holds. Otherwise, suppose without loss of generality that the right-hand side is not zero, so that equation 11 can be rewritten as
$\frac{\mu^{i}(E|h_{i}^{1})}{\mu^{i}(S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{2})|h_{i}^{1})}\cdot....\cdot\frac{\mu^{i}(S_{-i}(h_{i}^{n-1})\cap S_{-i}(h_{i}^{n})|h_{i}^{n})}{\mu^{i}(E|h_{i}^{n})}=1.$ (12)
Note that equation 12 holds if for every $s_{-i}\in E$ such that the ratio $\mu^{i}(s_{-i}|h_{i}^{1})/\mu^{i}(s_{-i}|h_{i}^{n})$ is not indeterminate,
$\frac{\mu^{i}(s_{-i}|h_{i}^{1})}{\mu^{i}(S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{2})|h_{i}^{1})}\cdot....\cdot\frac{\mu^{i}(S_{-i}(h_{i}^{n-1})\cap S_{-i}(h_{i}^{n})|h_{i}^{n})}{\mu^{i}(s_{-i}|h_{i}^{n})}=1,$ (13)
because this means that all the non-indeterminate ratios $\mu^{i}(s_{-i}|h_{i}^{1})/\mu^{i}(s_{-i}|h_{i}^{n})$ are identical, which implies that the ratio $\mu^{i}(E|h_{i}^{1})/\mu^{i}(E|h_{i}^{n})$ takes the same value too. We are going to show that for every $m=2,...,n$, there is $s_{-i}^{m}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$ such that
$\frac{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})|h_{i}^{m})}{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})|h_{i}^{m-1})}=\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m})}{\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})};$
then, we can rewrite equation 13 as
$\frac{\mu^{i}(s_{-i}|h_{i}^{1})}{\mu^{i}(s_{-i}^{2}|h_{i}^{1})}\cdot....\cdot\frac{\mu^{i}(s_{-i}^{n}|h_{i}^{n})}{\mu^{i}(s_{-i}|h_{i}^{n})}=1,$ (14)
which is true by complete consistency (cf. Remark 1). Let $s_{-i}^{m}$ be any $s_{-i}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$ such that $\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})>0$; one exists by the fact that the right-hand side of 11 is not zero. By complete consistency, for each $s_{-i}^{\prime}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$, we have
$\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m})}{\mu^{i}(s_{-i}^{\prime}|h_{i}^{m})}=\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})}{\mu^{i}(s_{-i}^{\prime}|h_{i}^{m-1})}.$
With some algebra, it is easy to show that then
$\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m})}{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})|h_{i}^{m})}=\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})}{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m-1})|h_{i}^{m-1})},$
as we wanted to show.
Only if part. By Remark 1 we can show that for each $s_{-i},\in S_{-i}$, for every generalized self-odds ratio, equation 14 holds. Without loss of generality, we can assume that the generalized self-odds ratio is not infinite (because an infinite one can be inverted and becomes zero), so that no denominator in equation 14 is zero. For each $m=2,...,n$, by consistency, equation 11 for $E=\left\{s_{-i}^{m}\right\}$ and $\left(h_{i}^{1},...,h_{i}^{n}\right)=(h_{i}^{m-1},h_{i}^{m})$ yields
$\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m})}{\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})}=\frac{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})|h_{i}^{m})}{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m-1})|h_{i}^{m-1})}.$
Then, substituing in 14, we obtain equation 13, which coincides with equation 12 for $E=\left\{s_{-i}\right\}$, and thus holds by consistency.
A.2 Completion of the proofs of the theorems
Proof of Theorem 1: $1\Rightarrow 2$
Fix a non-terminal contingency $h$ and let $h^{1},...,h^{n}$ be the next contingencies. Let $\alpha^{1},...,\alpha^{n}$ denote the weights that verify the definition of Bayes-plausibility. We construct probability distributions $(p(h^{m}|h,s))_{m=1}^{n}$ of reaching each $h^{m}$ given $h$ and each $s\in S(h)$. Fix $m\in\left\{1,...,n\right\}$. If $\alpha^{m}=0$, let $p(h^{m}|h,s)=0$ for all $s\in\mathcal{S}$. Suppose now that $\alpha^{m}>0$. For each $s\in\mathcal{S}$, let
$p(h^{m}|h,s)$ $=$ $\frac{\mu(s|h^{m})\alpha^{m}}{\mu(s|h)}\text{ \ \ if }s\in\mathrm{supp}\mu(\cdot|h),$
$p(h^{m}|h,s)$ $=$ $\alpha^{m}\text{ \ \ \ \ \ \ \ \ \ if }s\not\in\mathrm{supp}\mu(\cdot|h).$
For each $s\not\in\mathrm{supp}\mu(\cdot|h)$, we have
$\sum_{m=1}^{n}p(h^{m}|h,s)=\sum_{m=1}^{n}\alpha^{m}=1.$
For each $s\in\mathrm{supp}\mu(\cdot|h)$, we have
$\sum_{m=1}^{n}p(h^{m}|h,s)=\frac{\sum_{m=1}^{n}\mu(s|h^{m})\alpha^{m}}{\mu(s|h)}=1,$
where the last equality follows from Bayes-plausibility.
Now fix $m\in\left\{1,...,n\right\}$. If $\alpha^{m}=0$, then
$\sum_{s\in S(h)}p(h^{m}|h,s)=0.$
If $\alpha^{m}>0$, by Bayes-plausibility
$\mathrm{supp}\mu(\cdot|h^{m})\subseteq\mathrm{supp}\mu(\cdot|h).$ (15)
So we have
$\frac{p(h^{m}|h,s)\mu(s|h)}{\sum_{s^{\prime}\in S(h)}p(h^{m}|h,s^{\prime})\mu(s^{\prime}|h)}=\frac{\frac{\mu(s|h^{m})\alpha^{m}}{\mu(s|h)}\mu(s|h)}{\alpha^{m}\sum_{s^{\prime}\in\mathrm{supp}\mu(\cdot|h)}\mu(s|h^{m})}=\mu(s|h^{m}),\qquad\forall s\in\mathcal{S}$
where the last equality follows from (15). Thus, $\mu(\cdot|h^{m})$ can be derived with Bayes rule from $\mu(\cdot|h)$ using $(p(h^{m}|h,s))_{s\in S(h)}$ whenever possible.
Proof of Theorem 2: $3\Rightarrow 1$
We prove the counterpositive. Suppose that there exist two different GORs between some $s$ and $s^{\prime}$. Then, inverting all the DORs in the second GOR and multiplying by the first GOR, [^15] Since the two GORs are different, they cannot be both infinite or both zero, therefore after inverting one of the two their product is not indeterminate. we obtain a new GOR between $s$ and itself, which is not $1$; write it as
$r:=o(s^{0},s^{1}|h^{1})\cdot o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-1},s^{n}|h^{n}),$
where $n\geq 2$, $s^{0}=s^{n}=s$, and for each $k=1,...,n$,
$o(s^{k-1},s^{k}|h^{k})=\frac{p(h^{k}|s^{k-1})}{p(h^{k}|s^{k})}\frac{\mu(s^{k-1}|h^{k})}{\mu(s^{k}|h^{k})}.$ (16)
Wwithout loss of generality, we can suppose that:
$s^{0},...,s^{n-1}$ are all distinct, because if $s^{j}=s^{k}$, either $r^{\prime}=o(s^{j},s^{j+1}|h^{j})\cdot...\cdot o(s^{k-1},s|h^{k})$, or $r^{\prime\prime}=r/r^{\prime}$ is a new GOR between $s$ and itself that is not $1$.
$r<1$, because if a GOR is larger than $1$, inverting all the DORs we obtain a GOR smaller than $1$ (still between $s$ and itself).
With this, we construct a Dutch Book that the agent is willing to accept. For every $h\in\mathcal{H}\backslash\left\{h^{1},...,h^{n}\right\}$, let $g(\widetilde{s}|h)=0$ for every $\widetilde{s}\in\mathcal{S}$, so that $i$ is willing to accept $g(\cdot|h)$. Fix $\varepsilon>0$. Let $g(\cdot|h^{1})$ be the following gamble on $S(h^{1})$:
$g(s^{0}|h^{1})$ $=$ $-1,$
$g(s^{1}|h^{1})$ $=$ $-g(s^{0}|h^{1})\frac{\mu(s^{0}|h^{1})}{\mu(s^{1}|h^{1})}+\varepsilon=\frac{\mu(s^{0}|h^{1})}{\mu(s^{1}|h^{1})}+\varepsilon,$ (17)
and $g(s^{\prime}|h^{1})=0$ for every $s^{\prime}\not=s^{0},s^{1}$. Recursively, for each $m=2,...,n$, let $g(\cdot|h^{m})$ be the following gamble on $S(h^{m})$:
$g(s^{m-1}|h^{m})$ $=$ $-g(s^{m-1}|h^{m-1})\frac{p(h^{m-1}|s^{m-1})}{p(h^{m}|s^{m-1})}-\varepsilon,$ (18)
$g(s^{m}|h^{m})$ $=$ $-g(s^{m-1}|h^{m})\frac{\mu(s^{m-1}|h^{m})}{\mu(s^{m}|h^{m})}+\varepsilon,$ (19)
and $g(s^{\prime}|h^{m})=0$ for every $s^{\prime}\not=s^{m},s^{m-1}$. For each $m=1,...,n$, the agent is willing to accept $g(\cdot|h^{m})$ because
$\mu(s^{m-1}|h^{m})\cdot g(s^{m-1}|h^{m})+\mu(s^{m}|h^{m})\cdot g(s^{m}|h^{m})$
$=$ $\mu(s^{m-1}|h^{m})\cdot g(s^{m-1}|h^{m})-g(s^{m-1}|h^{m})\cdot\mu(s^{m-1}|h^{m})+\varepsilon\cdot\mu(s^{m}|h^{m})$
$=$ $\varepsilon\cdot\mu(s^{m}|h^{m})>0\text{,}$
where $\mu(s^{m}|h^{m})>0$ because $r\not=\infty$. Thus, $i$ is willing to accept $g=(g(\cdot|h))_{h\in\mathcal{H}}$. If $m\not=n$, since $s^{m}\not=s^{m^{\prime}}$ for every $m^{\prime}\not=m$, we also have
$\sum_{h\in\mathcal{H}}p(h|s^{m})g(s^{m}|h)$ $=$ $p(h^{m}|s^{m})g(s^{m}|h^{m})+p(h^{m+1}|s^{m})g(s^{m}|h^{m+1})$
$=$ $p(h^{m}|s^{m})g(s^{m}|h^{m})-p(h^{m+1}|s^{m})g(s^{m}|h^{m})\frac{p(h^{m}|s^{m})}{p(h^{m+1}|s^{m})}-p(h^{m+1}|s^{m})\varepsilon$
$=$ $-p(h^{m+1}|s^{m})\varepsilon<0,$
where $p(h^{m+1}|s^{m})>0$ because $s^{m}\in S(h^{m+1})$ (DORs at a contingency are only defined between consistent states). Thus, $s^{m}$ verifies the strict inequality condition of a Dutch book. Every $\widetilde{s}\in\mathcal{S}\backslash\left\{s^{1},...,s^{n}\right\}$ trivially satisfies the weak inequality condition. Finally, we need to show that, for some $\varepsilon>0$, $s$ satisfies the weak inequality condition as well, so that $g$ is a Dutch book. Note that, for $\varepsilon=0$,
$\sum_{h\in\mathcal{H}}p(h|s)g(s|h)$
$=$ $p(h^{1}|s)g(s|h^{1})+p(h^{n}|s)g(s|h^{n})$
$=$ $-p(h^{1}|s)-p(h^{n}|s)g(s^{n-1}|h^{n})\frac{\mu(s^{n-1}|h^{n})}{\mu(s|h^{n})}$
$=$ $-p(h^{1}|s)+p(h^{n}|s)g(s^{n-1}|h^{n-1})\frac{p(h^{n-1}|s^{n-1})}{p(h^{n}|s^{n-1})}\frac{\mu(s^{n-1}|h^{n})}{\mu(s|h^{n})}$
$=$ $-p(h^{1}|s)+p(h^{n-1}|s^{n-1})g(s^{n-1}|h^{n-1})o(s^{n-1},s|h^{n})$
$=$ $-p(h^{1}|s)-p(h^{n-1}|s^{n-1})g(s^{n-2}|h^{n-1})\frac{\mu(s^{n-2}|h^{n-1})}{\mu(s^{n-1}|h^{n-1})}\cdot o(s^{n-1},s|h^{n})$
$=$ $-p(h^{1}|s)+p(h^{n-1}|s^{n-1})g(s^{n-2}|h^{n-2})\frac{p(h^{n-2}|s^{n-2})}{p(h^{n-1}|s^{n-2})}\frac{\mu(s^{n-2}|h^{n-1})}{\mu(s^{n-1}|h^{n-1})}\cdot o(s^{n-1},s|h^{n})$
$=$ $-p(h^{1}|s)+p(h^{n-2}|s^{n-2})g(s^{n-2}|h^{n-2})o(s^{n-2},s^{n-1}|h^{n-1})\cdot o(s^{n-1},s|h^{n})$
$=$ $-p(h^{1}|s)-p(h^{1}|s^{1})g(s^{0}|h^{1})\frac{\mu(s|h^{1})}{\mu(s^{1}|h^{1})}\cdot o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-2},s^{n-1}|h^{n-1})\cdot o(s^{n-1},s|h)$
$=$ $-p(h^{1}|s)+p(h^{1}|s)r<0\text{,}$
where the second equality uses (19), the third (18), the fourth (16), and so on. By continuity, the strict inequality is preserved for sufficiently small $\varepsilon>0$. $\blacksquare$